(12) Solve the pair of simultaneous equations below:-

4a − 3b = 18

a + 2b = -1


(13) The molar heat capacity of a solid compound is given by c = a + bt


Given that when and that when c = 10 when t = 50 and that c = 13 when t = 80 Find the values of a and b.


(ALL WORKING OUTS TO BE SHOWN)


1
Expert's answer
2021-04-15T07:20:24-0400

(12)


"\begin{matrix}
 4a-3b=18 \\
a + 2b = -1
\end{matrix}"


"\begin{matrix}
 4a-3b-4a-8b=18-(-4) \\
a = -1-2b
\end{matrix}"

"\begin{matrix}
 -11b=22 \\
a = -1-2b
\end{matrix}"

"\begin{matrix}
 a=3 \\
b = -2
\end{matrix}"

(13)


"\begin{matrix}
 a+50b=10 \\
a + 80b = 13
\end{matrix}"


"\begin{matrix}
 a+80b-a-50b=13-10 \\
a = 10-50b
\end{matrix}"

"\begin{matrix}
 30b=3 \\
a =10-50b
\end{matrix}"

"\begin{matrix}
 a=5 \\
b = 0.1
\end{matrix}"